# The Silent Assistant: NoiseQuery as Implicit Guidance for Goal-Driven Image Generation

### v2

@@ Abstract
In this work, we introduce NoiseQuery as a novel method for enhanced noise initialization in versatile goal-driven text-to-image (T2I) generation. Specifically, we propose to leverage an aligned Gaussian noise as implicit guidance to complement explicit user-defined inputs, such as text prompts, for better generation quality and controllability. Unlike existing noise optimization methods designed for specific models, our approach is grounded in a fundamental examination of the generic finite-step noise scheduler design in diffusion formulation, allowing better generalization across different diffusion-based architectures in a tuning-free manner. This model-agnostic nature allows us to construct a reusable noise library compatible with multiple T2I models and enhancement techniques, serving as a foundational layer for more effective generation. Extensive experiments demonstrate that NoiseQuery enables fine-grained control and yields significant performance boosts not only over high-level semantics but also over low-level visual attributes, which are typically difficult to specify through text alone, with seamless integration into current workflows with minimal computational overhead.
@@ 2 Related Work
Enhancing T2I with Additional Conditions. To address the limitations of text-only conditions in meeting diverse user needs, numerous works [56, 6, 39, 12, 58, 7, 52] have expanded T2I generation through multi-modal conditions. For instance, methods like ControlNet [56] introduce structural signals (e.g., edge and depth maps) to control spatial layouts, while IC-Light [57] adjusts the illumination effects. Meanwhile, customization generation approaches focus on learning user-provided concepts like identity [12, 39, 40, 49, 55], style [48, 11], and interaction [21]. In this paper, we highlight that the initial noise for diffusion sampling can also serve as a conditional resource. Furthermore, compared to efforts that focus primarily on high-level semantic control, we show that initial noise holds untapped potential for guiding various low-level visual properties (e.g., color, texture, and sharpness).
Initial Noise Optimization for Diffusion Sampling. A growing body of research has focused on optimizing the initial noise in diffusion models, driven by the consensus that it significantly impacts the generated output. These methods iteratively refine initial noise through backpropagation at inference, guided by different generative objectives, such as semantic alignment [9, 15, 35, 3, 46], rare concept generation [42] and image translation [14]. For instance, ReNO [9] optimizes noises using various human preference rewards like HPS v2 [51], to enhance image quality and align with user intent. However, these approaches rely on iterative gradient updates during test-time, which are computationally expensive and prone to numerical instability (e.g., gradient explosion). In this work, we observe that initial noise exhibits cross-model consistency (Fig. 1), making it an exploitable shared asset. Leveraging this property, we construct a reusable noise library that not only facilitates efficient noise selection but also helps existing noise optimization methods improve generation quality.
@@ 3.3 Dissect Initial Noise via Generation Posterior
Cross-model Consistency. We also observe another meaningful phenomenon: the impact of initial noise is universal and model-agnostic. As illustrated in Fig. 1, generative posteriors generated from the same noise remain highly similar across models with different architectures and parameters. For example, PixArt-$\alpha$ [4] is a Transformer-based T2I diffusion model, while Stable Diffusion [37] is UNet-based. More discussion of models and samplers can be found in Appendix A. This finding empowers us to create a large-scale, model-agnostic noise library, where each noise sample can be associated with diverse latent features and then shared across different models. The universality of noise-driven priors positions initial noise as a unified control resource for various applications.
@@ 3.4 NoiseQuery for Goal-Driven Generation
$\epsilon^{*}=\arg\max_{\epsilon_{i}\in\mathcal{N}_{set}}\{S(\mathcal{F}_{i},\mathcal{F}_{O})\mid\epsilon_{i}\in\mathcal{N}_{set}\}.\vspace{-1mm}$ (4)
@@ 4 Experiments
+ ReNO [9] + NoiseQuery 1.71 23.52 32.92 32.78 - - - - 23.56 s
PixArt-$\alpha$ Base Model 0.70 22.08 28.27 30.83 0.78 22.24 29.33 31.48 4.327 s
@@ 4.1 Experimental Setup
Model Zoo. We conduct experiments on six popular diffusion generation models, including Stable diffusion (SD) series [37] (1.4, 1.5, 2.0, 2.1), PixArt-$\alpha$ [4], and SD-Turbo [43]. Stable Diffusion is a Latent Diffusion Model with UNet [38] backbone. By default, we use 50 DDIM denoising steps with a CFG scale of 7.0. PixArt-$\alpha$ is a Transformer Latent Diffusion Model, and SD-Turbo is a distillation model based on Stable Diffusion 2.1, generating high-quality images even with only 1-2 sampling steps without CFG. For PixArt-$\alpha$, we use the default CFG scale of 4.5.
@@ 4.2 Improvement on High-level Semantics
Quantitative Analysis. In Sec. 4, we evaluate our method on standard text-to-image (T2I) benchmarks, demonstrating consistent performance gains across all base models by offering better noise-text alignment. NoiseQuery also serves as a universal foundation layer, complementing T2I enhancement techniques like Diffusion-DPO [47], CFG++ [8], ReNO [9], and LaVi-Bridge [59]. Notably, these methods face inherent limitations: model-specific dependencies (e.g., Diffusion-DPO and LaVi-Bridge require retraining for new models; ReNO is restricted to one-step models due to heavy computation, excluding common multi-step models like SD) and high latency. In contrast, our method seamlessly integrates with them, enhancing their performance while adding near-zero time costs. We also present qualitative results in Fig. 5 to illustrate its effectiveness in assisting text-to-image synthesis. Additionally, we show the zero-shot transferability of NoiseQuery across various generative models in Sec. B.1. This model-agnostic compatibility, combined with minimal computational cost, establishes NoiseQuery as a versatile and efficient enhancement tool for diverse pipelines.
@@ 4.3 Controllability on Low-Level Visual Properties
Figure 10: Unlike the original Stable Diffusion , which produces images with medium brightness, our NoiseQuery (offset) expands the range to include both very bright and very dark samples.
Noise Offset via Generative Posteriors. Beyond directly selecting noise, we propose a method to introduce subtle color shifts by adjusting generative posteriors, preserving semantic content while allowing controlled color variations (details in Sec. C.4). This helps address the limitation of SD [37] in generating very bright or dark images, caused by noise distribution discrepancies during training and inference [28]. As shown in Fig. 10, subtly adjusting the brightness of generative posteriors can shift the inherent tendency of the initial noise, enabling the generation of images with a wider range of brightness during inference.
@@ 5 Conclusion
Acknowledgements. Research reported in this was supported by an Amazon Research Award Spring 2024 to OR and YZ. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not reflect the views of Amazon.
@@
In Appendix A, we introduce the generative posterior concept, emphasizing its consistency across different models and samplers. Moreover, Appendix B provides a detailed analysis of high-level semantic generation results, exploring our zero-shot transferability and the effects of CFG scales and library size on performance. Lastly, implementation details for experiments on low-level visual properties are outlined in Appendix C, featuring the noise offset method for color tendency control.
@@ Appendix A Analysis of Generative Posteriors
To further investigate the role of samplers, we compare generative posteriors from the same initial noise across multiple models (Stable Diffusion [37] v1.4, v1.5, v2.0, v2.1, SD-turbo [43] and PixArt-$\alpha$ [4]) and samplers (DDIM [44], LMS, Heun, Euler, PNDM [30], UniPC [60], DPM2 [22], DPM++ 2M [31], and DPM++ 2M Karras[31, 22]). As illustrated in Fig. S2, the generated images display remarkable consistency across different models and sampling strategies. This universal consistency enables the creation of a model-agnostic noise library, where each noise sample is associated with specific latent features and can be seamlessly reused across models and samplers. These findings underline the versatility of leveraging initial noise in a wide range of generative tasks.
@@ B.3 Different CFG Scale Analysis
Noise Size ImageReward PickScore HPS v2 CLIPScore
NoiseQuery (BLIP) 10k 0.4233 21.7495 0.2676 0.3162
50k 0.4374 21.7570 0.2682 0.3149
100k 0.4428 21.7577 0.2682 0.3146
200k 0.4569 21.7612 0.2683 0.3151
NoiseQuery (CLIP) 10k 0.3718 21.7349 0.2659 0.3146
50k 0.3677 21.7259 0.2655 0.3173
100k 0.3702 21.7349 0.2658 0.3171
200k 0.3645 21.7326 0.2653 0.3176
Random Noise 10k 0.3649 21.7264 0.2659 0.3138
50k 0.3707 21.7287 0.2663 0.3140
100k 0.3633 21.7211 0.2658 0.3137
200k 0.3648 21.7162 0.2658 0.3140
Table S2: Performance comparison of different noise library size on MSCOCO dataset.
B.4 Noise Library Size Analysis
In this section, we explore the impact of the noise library size on the performance of our method. Specifically, we investigate how increasing the size of the noise library affects various performance metrics, such as image reward, Pick Score [23], HPS v2 [51], and CLIP score [36], across different types of initial noise (NoiseQuery using BLIP and CLIP embeddings, as well as random noise). The results are presented in Tab. S2, which compares these metrics for different noise library sizes (10k, 50k, 100k, and 200k) on the MSCOCO [29] dataset. For NoiseQuery (BLIP), increasing the library size generally improves image quality, with the best results achieved at 200k samples. This suggests that a larger library provides more diverse noise samples, enabling the model to generate more accurate and refined outputs. For NoiseQuery (CLIP), a library size of 10k is sufficient for optimal performance, with no significant gains from increasing the library size further.
Regardless of the noise library size, NoiseQuery consistently outperforms random noise across all metrics, confirming the efficacy of our method in selecting appropriate noise for image generation.
@@ C.1 Texture
Texture refers to the small-scale spatial patterns of intensity variation in an image, which can be used to characterize the surface properties of objects. To control texture in generated images, we utilize the Gray-Level Co-occurrence Matrix (GLCM) [16], a commonly used technique for quantifying texture in terms of statistical measures. The GLCM features to capture the spatial arrangement and frequency of pixel intensity changes, providing a texture profile for each image.
@@ C.4 Noise Offset for Controlled Color Shifting
In addition to directly selecting the initial noise based on color attributes, we also propose a method to subtly shift the color tendency of a certain noise while preserving semantic content. We show the pseudo-code in  Algorithm 2 Specifically, we first generate an unconditional image from the initial noise and then apply subtle color adjustments, such as changes in brightness or saturation. After the color changes are made, we reverse the adjusted image back into the noise space via DDIM inversion. This results in a modified noise that consistently carries the desired color shifts, ensuring uniform color variations across different prompts while maintaining the underlying semantic content.

### v3

@@ Abstract
In this work, we introduce NoiseQuery as a novel method for enhanced noise initialization in versatile goal-driven text-to-image (T2I) generation. Specifically, we propose to leverage an aligned Gaussian noise as implicit guidance to complement explicit user-defined inputs, such as text prompts, for better generation quality and controllability. Unlike existing noise optimization methods designed for specific models, our approach is grounded in a fundamental examination of the generic finite-step noise scheduler design in diffusion formulation, allowing better generalization across different diffusion-based architectures in a tuning-free manner. This model-agnostic nature allows us to construct a reusable noise library compatible with multiple T2I models and enhancement techniques, serving as a foundational layer for more effective generation. Extensive experiments demonstrate that NoiseQuery enables fine-grained control and yields significant performance boosts not only over high-level semantics but also over low-level visual attributes, which are typically difficult to specify through text alone, with seamless integration into current workflows with minimal computational overhead. Code is available at https://github.com/wangruoyu02/NoiseQuery.
@@ 2 Related Work
Enhancing T2I with Additional Conditions. To address the limitations of text-only conditions in meeting diverse user needs, numerous works [64, 8, 45, 15, 66, 9, 59, 5, 13, 40] have expanded T2I generation through multi-modal conditions. For instance, methods like ControlNet [64] introduce structural signals (e.g., edge and depth maps) to control spatial layouts, while IC-Light [65] adjusts the illumination effects. Meanwhile, customization generation approaches focus on learning user-provided concepts like identity [15, 45, 46, 55, 63], style [54, 14], and interaction [25]. In this paper, we highlight that the initial noise for diffusion sampling can also serve as a conditional resource. Furthermore, compared to efforts that focus primarily on high-level semantic control, we show that initial noise holds untapped potential for guiding various low-level visual properties (e.g., color, texture, and sharpness).
Initial Noise Optimization for Diffusion Sampling. A growing body of research has focused on optimizing the initial noise in diffusion models, driven by the consensus that it significantly impacts the generated output [61, 24, 56, 36, 1]. These methods iteratively refine initial noise through backpropagation at inference, guided by different generative objectives, such as semantic alignment [11, 18, 41, 4, 52], rare concept generation [48] and image translation [17]. For instance, ReNO [11] optimizes noises using various human preference rewards like HPS v2 [58], to enhance image quality and align with user intent. However, these approaches rely on iterative gradient updates during test-time, which are computationally expensive and prone to numerical instability (e.g., gradient explosion). In this work, we observe that initial noise exhibits cross-model consistency (Fig. 1), making it an exploitable shared asset. Leveraging this property, we construct a reusable noise library that not only facilitates efficient noise selection but also helps existing noise optimization methods improve generation quality.
@@ 3.3 Dissect Initial Noise via Generation Posterior
[…]
@@ 3.4 NoiseQuery for Goal-Driven Generation
$\epsilon^{*}=\arg\max_{\epsilon_{i}\in\mathcal{N}_{set}}\{S(\mathcal{F}_{i},\mathcal{F}_{O})\mid\epsilon_{i}\in\mathcal{N}_{set}\}.\vskip-2.84526pt$ (4)
@@ 4.1 Experimental Setup
[…]
@@ 4.2 Improvement on High-level Semantics
Quantitative Analysis. In Tab. 2, we evaluate our method on standard text-to-image (T2I) benchmarks, demonstrating consistent performance gains across all base models by offering better noise-text alignment. NoiseQuery also serves as a universal foundation layer, complementing T2I enhancement techniques like Diffusion-DPO [53], CFG++ [10], ReNO [11], and LaVi-Bridge [67]. Notably, these methods face inherent limitations: model-specific dependencies (e.g., Diffusion-DPO and LaVi-Bridge require retraining for new models; ReNO is restricted to one-step models due to heavy computation, excluding common multi-step models like SD) and high latency. In contrast, our method seamlessly integrates with them, enhancing their performance while adding near-zero time costs. We also present qualitative results in Fig. 5 to illustrate its effectiveness in assisting text-to-image synthesis. Additionally, we show the zero-shot transferability of NoiseQuery across various generative models in Sec. B.1. This model-agnostic compatibility, combined with minimal computational cost, establishes NoiseQuery as a versatile and efficient enhancement tool for diverse pipelines.
@@ 5 Conclusion
Limitations. Although NoiseQuery provides flexibility for aligning high-level semantics and controlling low-level attributes, it remains bounded by the base model’s ability, and may still underperform in extremely complex cases. Additionally, it has limitations in achieving fine-grained control, such as Canny edge control.
Acknowledgements
This work was partially supported by the National Natural Science Foundation of China under grant 62372341. This work was also partially supported by an Amazon Research Award 2023 to OR and YZ. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not reflect the views of Amazon.
@@ Acknowledgements
In Appendix A, we introduce the generative posterior concept, emphasizing its consistency across different models and samplers. Moreover, Appendix B provides a detailed analysis of high-level semantic generation results, exploring our zero-shot transferability, complex image generation, and the effects of CFG scales. Lastly, implementation details for experiments on low-level visual properties are outlined in Appendix C, along with the noise offset method for color tendency control.
@@ Appendix A Analysis of Generative Posteriors
[…]
@@ B.1 Zero-Shot Transferability
Additionally, NoiseQuery provides a goal-aligned initial noise while users can freely choose various schedulers for generation. As shown in Tab. S2, a library built with DDIM generalizes well to other schedulers.
@@
[…]
PDNM SD v2.1 / +NoiseQuery 0.10 / 0.26 21.31 / 21.42 24.71 / 25.23 30.91 / 31.67
DDIM SD v2.1 / +NoiseQuery 0.12 / 0.26 21.30 / 21.45 24.72 / 25.17 31.18 / 31.71
LMS SD v2.1 / +NoiseQuery 0.07 / 0.27 21.31 / 21.48 24.66 / 25.37 30.83 / 31.76
Heun SD v2.1 / +NoiseQuery 0.09 / 0.27 21.34 / 21.48 24.66 / 25.29 30.93 / 31.23
Euler SD v2.1 / +NoiseQuery 0.12 / 0.22 21.35 / 21.38 24.61 / 25.12 31.23 / 31.54
DPM2 SD v2.1 / +NoiseQuery 0.09 / 0.27 21.31 / 21.42 24.66 / 25.37 30.81 / 31.79
DPM++ 2M Karras SD v2.1 / +NoiseQuery 0.10 / 0.27 21.36 / 21.44 24.77 / 25.32 30.83 / 31.79
DPM++ 2M SDE Karras SD v2.1 / +NoiseQuery 0.23 / 0.28 21.42 / 21.59 23.32 / 25.68 31.20 / 31.23
Table S2: Zero-shot transferability from DDIM to various samplers on DrawBench [47].
@@ C.1 Texture
Texture refers to the small-scale spatial patterns of intensity variation in an image, which can be used to characterize the surface properties of objects. To control texture in generated images, we utilize the Gray-Level Co-occurrence Matrix (GLCM) [19], a commonly used technique for quantifying texture in terms of statistical measures. The GLCM features capture the spatial arrangement and frequency of pixel intensity changes, providing a texture profile for each image.
@@ C.2 Shape
Figure S4: Unlike the original Stable Diffusion, which produces images with medium brightness, our NoiseQuery (offset) expands the range to include both very bright and very dark samples.
@@ C.4 Noise Offset for Controlled Color Shifting
Beyond directly selecting noise, we also propose a method to introduce subtle color shifts by adjusting generative posteriors, preserving semantic content while allowing controlled color variations. We show the pseudo-code in  Algorithm 2 Specifically, we first generate an unconditional image from the initial noise and then apply subtle color adjustments, such as changes in brightness or saturation. After the color changes are made, we reverse the adjusted image back into the noise space via DDIM inversion. This results in a modified noise that consistently carries the desired color shifts, ensuring uniform color variations across different prompts while maintaining the underlying semantic content.
This helps address the limitation of SD [43] in generating very bright or dark images, caused by noise distribution discrepancies during training and inference [32]. As shown in Fig. S4, subtly adjusting the brightness of generative posteriors can shift the inherent tendency of the initial noise, enabling the generation of images with a wider range of brightness during inference. shifts the noise distribution, enabling a wider range of brightness in generated images.